OpenStudy (anonymous):

Simplify 14 to the 15th power over 14 to the 5th power. 13 110 143 1410

OpenStudy (anonymous):

@Adjax

OpenStudy (anonymous):

subtract the exponents

OpenStudy (anonymous):

\[\huge \frac{b^m}{b^n}=b^{m-n}\]

OpenStudy (anonymous):

\[\frac{ 14^{15} }{ 14^5 } = 14^{15-5} = 14^{10}\]

OpenStudy (mathmath333):



OpenStudy (anonymous):

=289254654976

OpenStudy (anonymous):

simplyfy that rt

OpenStudy (anonymous):

14 10 is the answer